# Reinforcement Learning with Elastic Time Steps

### v3

@@ VI Conclusions
To prevent reward explosion, we introduced an upper limit ($\alpha_{max}$) for the hyperparameter $\alpha_{m}$, ensuring convergence and stability, validated using a Lyapunov model. The boundedness analysis confirmed that $\alpha_{m}$ remains within the defined limits, maintaining the strength of the reward function and ensuring the robustness and reliability of the learning process.
@@
Building on MOSEAC’s success, our future work will focus on extending the principles of variable time step algorithms to other reinforcement learning frameworks, such as Hierarchical Reinforcement Learning (HRL). This extension aims to address the complexities of long-term planning tasks more efficiently, further improving the adaptability and performance of RL algorithms in diverse and challenging environments, which brings benefits for solving complex problems by RL using real robot systems.

### v4

@@ 6 Conclusions
To prevent reward explosion, we introduced an upper limit ($\alpha_{max}$) for the hyperparameter $\alpha_{m}$, ensuring convergence and stability, validated utilizing a Lyapunov model. The boundedness analysis confirmed that $\alpha_{m}$ remains within the defined limits, maintaining the strength of the reward function and ensuring the robustness and reliability of the learning process.
@@
Building on MOSEAC’s success, our future work will focus on extending the principles of variable time step algorithms to other RL frameworks, such as Hierarchical Reinforcement Learning (HRL). This extension aims to address the complexities of long-term planning tasks more efficiently, further improving the adaptability and performance of RL algorithms in diverse and challenging environments, which brings benefits for solving complex problems by RL utilizing real robot systems.